# PartDexTOG: Generating Dexterous Task-Oriented Grasping via Language-driven Part Analysis

### v1

@@ 1 Introduction
Task-Oriented Grasping (TOG), which requires selecting the optimal area of an object to grasp, such that the execution of the subsequent manipulation could be facilitated, has recently attracted extensive research attention [1; 2]. Existing works have explored achieving TOG through building knowledge graphs [1; 3], learning from videos with human actions [4; 5], and utilizing pre-trained foundation models [6; 7]. The generated versatile grasping enables robots to interact with objects in human-like manners [2; 8] and would boost a variety of downstream applications [9; 10; 11].
@@
In this work, we advocate generating dexterous TOG with the aid of part analysis. Parts are fundamental elements in understanding how objects can be effectively interacted with. On the other hand, recent works in Large Language Models (LLMs) have significantly advanced robotic manipulation, enabling zero-shot generalization in unseen scenarios [13; 14], thanks to the open-world knowledge reasoning capabilities of LLMs. By incorporating LLMs, part analysis can enhance robot grasping by providing the part-level functionality of open-world objects, eliminating the necessity for training on large-scale grasping datasets.
@@ 2.1 Task-oriented grasping
TOG requires detecting grasps considering the manipulation task. Early works employed analytical methods [15; 16; 17; 18; 19] to optimize task wrench space. These methods lack generality, limiting their applications. More recently, data-driven approaches have greatly boosted the area. For example, GCNGrasp [1] constructs knowledge graphs to filter grasping strategies. GraspGPT [6] and FoundationGrasp [7] introduce LLMs to extract open-world semantic information. LERF-TOGO [20] generates the grasp distribution on objects utilizing LERF [21]. SegGrasps [22] utilizes vision-language models and convex decomposition to achieve object segmentation for grasping.
Several recent works adopt dexterous hands to conduct TOG, bringing better flexibility [23; 24]. TaskDexGrasp [25] aligns task and grasp wrench space. DexGYSGrasp [26] utilizes CLIP to achieve the analysis of task semantics. DexFG-Net [27] is based on a dense shape correspondence. NL2Contact [28] generates controllable grasps by using a multi-stage diffusion model. DexTOG [29] uses the intermediate generation grasp as control conditions for generating the grasp. SemGrasp [30] uses multi-modal LLMs to achieve cross-modal alignment and integration. SayFuncGrasp [31] utilizes LLMs to evaluate grasping poses and parts. AffordDexGrasp [32] introduces new Generalizable-Instructive Affordance. TextGraspdiff [33] and G-DexGrasp [34] both utilize object parts. Although they share a similar concept of using object parts, our method distinguishes itself by not requiring a given grasping part, which is a more general solution for downstream applications.
@@ 2.2 LLMs for robotics
LLMs have recently gained significant attention in the field of robotics. The unparalleled capabilities in open-world understanding and reasoning enable LLMs to boost various robotics applications [35; 36; 37]. LLMs could interpret the instructions and generate sequential actions by leveraging high-level semantic knowledge about tasks [38; 39]. These methods show good zero-shot generalization in task understanding and high-quality real-world grounding to ensure the planned actions are feasible and contextually appropriate [40; 41; 42; 43]. Our method follows similar protocols of existing works. It explores leveraging LLMs to boost a specific task, i.e. TOG, through extracting the part functionalities from language. The proposed framework as well as the designed LLM prompts are expected to be useful for other relevant robotic tasks.
@@ 3.3 Dexterous grasp generation
[…]
@@
where $\text{PC}^{\text{Hand}}_{n}$ is the point cloud of the generated hand’s surface and $\text{PC}^{\text{Part}}_{n}$ is the point cloud of the part. $d(\cdot)$ is the L2 distance. $f[\cdot]$ returns $1$ if the input value is less than a threshold $\lambda$, and returns 00 otherwise. $|\text{PC}^{\text{Part}}_{n}|$ is the number of points in the part. We select the grasp with the highest $E_{n}$ as the most plausible grasp. A visualization of the selection is shown in Figure 4.
@@ 4.1 Dataset
We train and test PartDexTOG on the OakInk-shape dataset [49], the largest dataset of dexterous TOG, containing $50$k hand-object poses and models. The diversities of shapes and tasks allow comprehensive evaluations of the proposed method. Specifically, the training set and testing set contains grasps on $923$ and $176$ objects, respectively. The latter includes objects from untrained categories, facilitating the comprehensive evaluation of generality. This dataset is the most inclusive, diverse, and versatile one for dexterous TOG [57]. Evaluating on this dataset only is sufficient to demonstrate the advantageous performance of our method.
@@ 4.2 Evaluation Metrics
Moreover, to evaluate the semantic correctness, we employ the following metrics: 1) P-FID: The Fréchet Inception Distance between the PointNet++ feature of the predicted and ground-truth hand [22; 61]; 2) LLM score: A semantic score generated by the GPT-4v. We use the same setting as [22]; 3) Diversity: The variance of the grasp parameters. Higher variances represent more diversity [62]; 4) Perceptual Scores: We conduct user studies to evaluate the generated grasps. Each volunteer is asked to evaluate $500$ individual grasps, assigning perceptual scores between 00-$5$ in the following aspects: a) Semantic Consistency (SC) between the task and the grasp; b) Physical Plausibility (PP) of hand posture; c) Interaction Stability (IS) between hand and object; d) Correctness of Part selection (CP).
@@
Method Penetration Volume$(cm^{3})$$\downarrow$ Grasp Displace$(cm)$$\downarrow$ Penetration Volume$(cm^{3})$$\downarrow$ Grasp Displace$(cm)$$\downarrow$
@@ A.3 Discussion of Baseline Methods in Task-oriented Dexterous Grasping
Our method generates dexterous task-oriented grasping from a given task type and a 3D object. As discussed in the introduction and related work sections, few existing works address this challenging task. Several concurrent works study a similar problem. We did not compare with those methods due to the significant difference in problem settings [65; 66; 25; 32] or the unreleased source code/training data [66; 34; 28].

### v2

@@ 1 Introduction
Task-Oriented Grasping (TOG), which requires selecting the optimal area of an object to grasp, such that the execution of the subsequent manipulation could be facilitated, has recently attracted extensive research attention [40; 29]. Existing works have explored achieving TOG through building knowledge graphs [40; 2], learning from videos with human actions [43; 14], and utilizing pre-trained foundation models [52; 51]. The generated versatile grasping enables robots to interact with objects in human-like manners [29; 61] and would boost a variety of downstream applications [11; 8; 9].
@@
In this work, we advocate generating dexterous TOG with the aid of part analysis. Parts are fundamental elements in understanding how objects can be effectively interacted with. On the other hand, recent works in Large Language Models (LLMs) have significantly advanced robotic manipulation, enabling zero-shot generalization in unseen scenarios [60; 3], thanks to the open-world knowledge reasoning capabilities of LLMs. By incorporating LLMs, part analysis can enhance robot grasping by providing the part-level functionality of open-world objects, eliminating the necessity for training on large-scale grasping datasets.
@@ 2.1 Task-oriented grasping
TOG requires detecting grasps considering the manipulation task. Early works employed analytical methods [4; 17; 34; 15; 16] to optimize task wrench space. These methods lack generality, limiting their applications. More recently, data-driven approaches have greatly boosted the area. For example, GCNGrasp [40] constructs knowledge graphs to filter grasping strategies. GraspGPT [52] and FoundationGrasp [51] introduce LLMs to extract open-world semantic information. LERF-TOGO [45] generates the grasp distribution on objects utilizing LERF [27]. SegGrasps [30] utilizes vision-language models and convex decomposition to achieve object segmentation for grasping.
Several recent works adopt dexterous hands to conduct TOG, bringing better flexibility [39; 36]. TaskDexGrasp [7] aligns task and grasp wrench space. DexGYSGrasp [57] utilizes CLIP to achieve the analysis of task semantics. DexFG-Net [56] is based on a dense shape correspondence. NL2Contact [65] generates controllable grasps by using a multi-stage diffusion model. DexTOG [64] uses the intermediate generation grasp as control conditions for generating the grasp. SemGrasp [31] uses multi-modal LLMs to achieve cross-modal alignment and integration. SayFuncGrasp [35] utilizes LLMs to evaluate grasping poses and parts. AffordDexGrasp [58] introduces new Generalizable-Instructive Affordance. TextGraspdiff [6] and G-DexGrasp [24] both utilize object parts. Although they share a similar concept of using object parts, our method distinguishes itself by not requiring a given grasping part, which is a more general solution for downstream applications.
@@ 2.2 LLMs for robotics
LLMs have recently gained significant attention in the field of robotics. The unparalleled capabilities in open-world understanding and reasoning enable LLMs to boost various robotics applications [48; 19; 5]. LLMs could interpret the instructions and generate sequential actions by leveraging high-level semantic knowledge about tasks [63; 54]. These methods show good zero-shot generalization in task understanding and high-quality real-world grounding to ensure the planned actions are feasible and contextually appropriate [28; 20; 22; 23]. Our method follows similar protocols of existing works. It explores leveraging LLMs to boost a specific task, i.e. TOG, through extracting the part functionalities from language. The proposed framework as well as the designed LLM prompts are expected to be useful for other relevant robotic tasks.
@@ 3.3 Dexterous grasp generation
$F^{\text{Cat}}=\operatorname{BERT}\left(\emph{D[object\_category , task\_type] }\right),\quad F^{\text{Part}}_{n}=\operatorname{BERT}\left(\emph{D[part\_label\_n] }\right).$ (1)
@@
where $\text{PC}^{\text{Hand}}_{n}$ is the point cloud of the generated hand’s surface and $\text{PC}^{\text{Part}}_{n}$ is the point cloud of the part. $d(\cdot)$ is the L2 distance. $f[\cdot]$ returns $1$ if the input value is less than a threshold $\lambda$, and returns 0 otherwise. $|\text{PC}^{\text{Part}}_{n}|$ is the number of points in the part. We select the grasp with the highest $E_{n}$ as the most plausible grasp. A visualization of the selection is shown in Figure 4.
@@ 4.1 Dataset
We train and test PartDexTOG on the OakInk-shape dataset [62], the largest dataset of dexterous TOG, containing 5050k hand-object poses and models. The diversities of shapes and tasks allow comprehensive evaluations of the proposed method. Specifically, the training set and testing set contains grasps on $923$ and $176$ objects, respectively. The latter includes objects from untrained categories, facilitating the comprehensive evaluation of generality. This dataset is the most inclusive, diverse, and versatile one for dexterous TOG [59]. Evaluating on this dataset only is sufficient to demonstrate the advantageous performance of our method.
@@ 4.2 Evaluation Metrics
Moreover, to evaluate the semantic correctness, we employ the following metrics: 1) P-FID: The Fréchet Inception Distance between the PointNet++ feature of the predicted and ground-truth hand [30; 42]; 2) LLM score: A semantic score generated by the GPT-4v. We use the same setting as [30]; 3) Diversity: The variance of the grasp parameters. Higher variances represent more diversity [55]; 4) Perceptual Scores: We conduct user studies to evaluate the generated grasps. Each volunteer is asked to evaluate $500$ individual grasps, assigning perceptual scores between 0-55 in the following aspects: a) Semantic Consistency (SC) between the task and the grasp; b) Physical Plausibility (PP) of hand posture; c) Interaction Stability (IS) between hand and object; d) Correctness of Part selection (CP).
@@
Method $downarrow Penetration Volume(cm3)(cm^{3})\downarrow Grasp Displace(cm)(cm)\downarrow$
@@ A.3 Discussion of Baseline Methods in Task-oriented Dexterous Grasping
Our method generates dexterous task-oriented grasping from a given task type and a 3D object. As discussed in the introduction and related work sections, few existing works address this challenging task. Several concurrent works study a similar problem. We did not compare with those methods due to the significant difference in problem settings [13; 66; 7; 58] or the unreleased source code/training data [66; 24; 65].